# Claim Verification in the Age of Large Language Models: A Survey

### v1

@@ Introduction
Although recent studies show the success of pre-trained LLMs on zero- or few-shot tasks, they often fail to verify real-world claims given their limited internal knowledge. The success of fine-tuning has motivated recent work on claim verification (Chen et al. 2022b; Pan et al. 2021; Zeng and Zubiaga 2024). The work in Chen et al. (2022b) leverages a language model to fine-tune over an external corpus for retrieving passage titles and evidence sentences using constrained beam search. The results showed improved performance against the traditional fact verification pipeline over the FEVER dataset. Other work demonstrates that using GPT models to generate synthetic training data improves the performance of LLMs on various tasks such as fact checking (Tang, Laban, and Durrett 2024) and claim matching (Choi and Ferrara 2024).
Authors in Pan et al. (2021) develop a pipeline for creating a fact verification dataset and fine-tuning a language model by leveraging passages from Wikipedia to generate QA pairs related to claim veracity. The authors show that it can improve state-of-the-art language models in zero-shot settings. The work in (Zeng and Zubiaga 2024) shows that using unlabelled pairwise data to increase the alignment between claim-evidence pairs shows significant improvement of LLM performance in few-shot claim verification tasks. In addition, other recent works leverage techniques such as reinforcement learning to fine-tune models for improving the veracity of claims and supporting evidence Zhang and Gao (2024); Huang et al. (2024). A document-level and question-level retrieval policy is proposed by Zhang and Gao (2024), where the top-k and top-1 documents for the document and question-level policy respectively are used as input to a scoring function for label prediction during training. This approach outperforms retrieval and prompting approaches. Chiang et al. (2024) fine-tune LLMs for their multi-stage fact verification approach. They fine-tune a model to generate answers based on claim-evidence pairs and a set of questions, whereas another model is fine-tuned to verify the claim based on the claim-evidence and question-answer pairs. Authors in Zhu et al. (2023) fine-tune a generation model to generate counterfactuals to train the fact verification model’s performance on out-of-domain claims.
@@
The recent success in the performance of pre-trained LLMs in zero- and few-shot settings is largely attributed to its in-context learning (ICL) abilities (Kojima et al. 2022; Brown et al. 2020). In the scope of claim verification, popular ICL techniques with strong zero- and few-shot performance include chain-of-thought (CoT) reasoning (Wei et al. 2022). The work in Zhao et al. (2024) develops a multi-stage verification pipeline for claim verification based on claim decomposition and self-reflection. An LLM-based verifier module is created using instruction prompting to generate a reasoning analysis among all sub-claims created by the decomposer module. The results suggest that using zero-shot prompting techniques provides better performance in multi-hop claim verification domains such as HOVER and FEVEROUS compared to few-shot prompting and fine-tuning methods. Authors in Kanaani (2024) enable LLMs to generate reasons over retrieved evidence in claim verification using few-shot ICL and the STaR CoT technique inspired by the work of Zelikman et al. (2022). Similar work has leveraged CoT techniques for the purposes of effectively verifying complex claims using reasoning steps (Yao et al. 2023; Ni et al. 2024). On the other hand, HiSS (Zhang and Gao 2023) demonstrates that using prompting techniques for few-shot learning and claim decomposition can substantially improve the performance of CoT models for complex news claim verification. Li et al. (2023b) leverage the ICL capability of LLMs to perform multiple tasks simultaneously. Their approach outperforms or achieves comparable task performance in a zero-shot setting on claim verification datasets.
LLM Generation Strategies
@@
While the majority of claim verification systems predict veracity labels based on the concatenation of the input claim and the set of evidence sentences (Pradeep et al. 2021), recent work has proposed alternate strategies for determining veracity labels as well as selecting/generating evidence pieces. The work in Cao et al. (2024) develops, SERIf, a claim verification pipeline that features an inference module that predicts the veracity label of scientific news articles based on a two-step summarization (i.e., ‘Extractive - Abstractive‘) and evidence retrieval technique. Each summary-evidence pair is fed into the LLM and it produces a binary label, indicating whether the news article is reliable (supported) or unreliable (refuted). Authors in Wadden et al. (2022b) leverage the Longformer transformer model (Beltagy, Peters, and Cohan 2020) that uses a shared encoding over the claim and document abstracts for rationale identification and claim label prediction.
@@
The work in Li et al. (2024a) develops an algorithm that selects the minimal evidence group (MEG) within a set of retrieved candidate documents. The algorithm aims to minimize the redundancy while also selecting the most relevant piece of evidence to prompt the language model. Authors in Chen et al. (2022b) create an LLM claim verification framework in which the authors use the BART model to encode all candidate sentences from the most relevant retrieved documents. The BART decoder serves as an evidence decoder to predict the $g$-th evidence sentence via generation conditioned on the top $k$ retrieved documents and the input claim. Lee et al. (2022) developed a variant of nucleus sampling called, factual-nucleus sampling, in which the top-$p$ sampling pool is selected as a set of sub-words whose cumulative probability exceeds $p$. The authors show that factual-nucleus sampling can improve evidence and label generation without claim verification datasets such as FEVER. Kao and Yen (2024) propose a multi-stage approach, where the evidence sentences are retrieved from articles related to a claim, and arguments are generated by aggregating and reconstructing the evidence. The arguments are refined and passed to a LLM to generate a verification label. Other approaches leverage LLMs reasoning capabilities to generate veracity labels and factual evidence (Cheng, Tan, and Lu 2024; Jafari and Allan 2024; Li et al. 2024b; Fang et al. 2024; Pan et al. 2023a).
Dataset # Instances Source # Labels Reference
AVeriTeC 4,568 Google FactCheck Claim Search API 4 Schlichtkrull, Guo, and Vlachos (2024)
BINGCHECK 3,840 Microsoft Copilot 4 Li et al. (2023c)
Check-COVID 1,504 News websites, CORD-19 3 Wang et al. (2023a)
CLAIMDECOMP 1,000 PolitiFact 6 Chen et al. (2022a)
CLIMATE-FEVER 1,535 Google Search, Wikipedia 3 Diggelmann et al. (2020)
CoVERT 212 Manually created 3 Liu et al. (2024)
COVID-Fact 4,086 Reddit 2 Saakyan, Chakrabarty, and Muresan (2021)
EX-FEVER 60,000 Wikipedia 3 Ma et al. (2024)
ExpertQA 2177 LLM, Google Search 5 Malaviya et al. (2024)
FACTIFY-5WQA 391,041 Public Fact-Verification Datasets 3 Rani et al. (2023)
FactCheck-Bench 678 LLM 4 Wang et al. (2023b)
FactCheckQA 20,871 PolitiFact 3 Bashlovkina et al. (2023)
FACTUALITYPROMPTS 16,000 Wikipedia, LLM 2 Lee et al. (2022)
FAVIQ 188,000 Wikipedia, Existing Datasets 2 Park et al. (2022)
FEVER 185,445 Wikipedia 3 Thorne et al. (2018a)
FEVEROUS 87,026 Wikipedia 3 Aly et al. (2021)
FoolMeTwice 12,968 Wikipedia 2 Eisenschlos et al. (2021)
HealthVer 14,330 Bing Web Search API 3 Sarrouti et al. (2021)
HOVER 26,171 Wikipedia 2 Jiang et al. (2020)
LIAR 12,500 PolitiFact 6 Wang (2017)
MultiFC 34,918 Fact-checking Websites 2-40 Augenstein et al. (2019)
PolitiFact 21,152 PolitiFact 6 Misra (2022)
QuanTemp 15,514 Google Fact Check Tool API 3 Venktesh et al. (2024)
RAWFC 2,012 Snopes 6 Yang et al. (2022)
SciFact 1,400 S2ORC 2 Wadden et al. (2020)
SciFact-Open 279 S2ORC 3 Wadden et al. (2022a)
SciNews 2,400 Wiki Fake News, LLM, Existing Datasets 2 Cao et al. (2024)
SCITAB 1,225 SciGen Dataset (arXiv) 3 Lu et al. (2023)
VitaminC 400,000 Wikipedia 3 Schuster, Fisch, and Barzilay (2021)
WiCE 1,967 Wikipedia 3 Kamoi et al. (2023)
XClaimCheck 16,177 PolitiFact 5 Kao and Yen (2024)
Table 1: Available datasets along with the number of instances, data sources, and references.
@@
Recent studies investigate explainable approaches to improve LLM-based claim verification systems (Wang and Shu 2023; Dammu et al. 2024; Si et al. 2023). Authors in Wang and Shu (2023) present an explainable claim verification framework named FOLK, that leverages the explanation capabilities of LLMs when verifying a claim and justifies the prediction through a summary of its decision process. The work in Dammu et al. (2024) proposes a knowledge graph (KG)-based approach for text verification and evidence attribution. An objective function is used to fine-tune LLMs on evidence attribution based on the input text and retrieved triplets from KG, inducing explanations on claim predictions. While explainable techniques can aid humans in fact-checking, LLMs are prone to incorrect explanations due to hallucinations, causing them to be unreliable in certain claim verification scenarios (Si et al. 2023). Ma et al. (2024) prompt an LLM in a few-shot setting to generate a concise summary of the gathered evidence documents and the input claim. This summary serves as an explanation for the verified claim. Other works in claim verification leverage reasoning techniques such as chain-of-thought (CoT) for enabling the LLM to be interpretable in its decision-making process when verifying claims (Yao et al. 2023; Pan et al. 2023a; Zhao et al. 2024; Kanaani 2024; Ni et al. 2024; Quelle and Bovet 2024; Fang et al. 2024). The authors in Pan et al. (2023a) and Fang et al. (2024) leverage the reasoning ability of the LLMs to generate explanations by using question-guided reasoning and minimizing the inherent model biases respectively.
Evaluation and Benchmarking
Metrics
@@
The F1 score is the most commonly used metric to measure the performance of automatic claim verification systems. Other metrics like Precision, Recall, and Accuracy are also used to evaluate the system performance. Katranidis and Barany (2024) used the error rate between the human and automated fact-verification system to measure the verification accuracy. However, these metrics consider a single pipeline component to evaluate the system’s overall performance. Hence, Thorne et al. (2018a) introduced the FEVER score, a metric that uses both the verification accuracy as well as the evidence retrieval accuracy to compute overall system performance. While these metrics are valuable for assessing the performance of the classification tasks, they are inadequate for evaluating the performance of the non-classification components of the pipeline. Hence, metrics like Recall@k are used to measure the performance of the retrieval task (Pan et al. 2023b; Pradeep et al. 2021), while BLEU and METEOR are used to evaluate the quality of explanations or generated questions and answers. Schlichtkrull, Guo, and Vlachos (2024) propose a new evaluation metric AVeriTeC score that uses METEOR and accuracy metrics, for question-answer-based veracity label prediction systems. Other metrics like Mean Absolute Error (MAE), Expected Calibration Error (ECE), Area Under ROC Curve (AUC-ROC), and Pearson’s Correlation are also used. Most of these metrics are inadequate to evaluate the factual accuracy of the LLM-generated text. Several metrics like FactScore (Min et al. 2023), SAFE (Wei et al. 2024), and VERISCORE (Song, Kim, and Iyyer 2024) are introduced to evaluate the factual accuracy of the LLM-generated text. While these metrics evaluate the accuracy, other metrics and frameworks to evaluate the factual errors in generated text (Lee et al. 2022; Chern et al. 2023) as well as their alignment (Zha et al. 2023) and entailment (Lee et al. 2022) considering the factuality have been proposed.
Datasets
@@
The fundamental resource for training and evaluating claim verification systems is datasets containing annotated texts. As most research in this area deals with English data, we collect information about all publicly available English datasets used in the papers discussed in this survey and present them in Table 1.
@@
Several general-domain datasets have been released over the years. Different online data sources and websites are used to extract facts and claims including Wikipedia (Thorne et al. 2018a; Jiang et al. 2020; Diggelmann et al. 2020; Eisenschlos et al. 2021; Schuster, Fisch, and Barzilay 2021; Kamoi et al. 2023), due to the extensive amount of information available spanning various topics and domains, and online fact-checking websites like PolitiFact (Wang 2017; Augenstein et al. 2019; Kao and Yen 2024). Factually incorrect claims are shared through social media channels like X, Reddit, etc. Saakyan, Chakrabarty, and Muresan (2021) introduced the COVID-Fact dataset consisting of claims extracted from Reddit posts. Several datasets for scientific fact verification have also been introduced over the years (Wadden et al. 2020, 2022a; Lu et al. 2023). Most fact-verification datasets rely on unstructured textual evidence. Hence, a few datasets with structured data sources as evidence have been introduced (Aly et al. 2021; Lu et al. 2023). While a large number of datasets are focused on only veracity labels, some datasets were deloped to address explainability as well (Chen et al. 2022a; Yang et al. 2022; Ma et al. 2024; Rani et al. 2023; Schlichtkrull, Guo, and Vlachos 2024).
@@
Most of the claim-verification datasets include claims extracted from available information sources. LLMs are widely used to generate content, given their superior text-generation capabilities. These models often generate factually incorrect information. Efforts to identify such non-factual text have been undertaken. However, existing claim-verification datasets can be inadequate for this task due to the linguistic variations between the human- and LLM-generated text. To overcome this issue, LLM-generated claim verification datasets have been introduced (Li et al. 2023c; Cao et al. 2024). While these datasets are used to evaluate the automatic claim verification systems, there is a need to evaluate the factual accuracy of the LLM-generated text. Hence, a few datasets to evaluate the LLM generation’s factual accuracy have been introduced (Lee et al. 2022; Wang et al. 2023b; Malaviya et al. 2024).
Shared Tasks
@@
Shared tasks are competitions where participating teams develop systems to solve a task using a common benchmark dataset. There have been multiple shared tasks on the claim and fact-checking organized over the years like the Fact Extraction and Verification (FEVER) (Thorne et al. 2018b), TabFact [^4] https://competitions.codalab.org/competitions/21611, CLEF 2020 CheckThat! (Barrón-Cedeno et al. 2020), SCIVER (Wadden and Lo 2021), SEM-TAB-FACT (Wang et al. 2021), FACTIFY-5WQA [^5] https://defactify.com/factify3.html, and more recently AVeriTeC [^6] https://fever.ai/task.html. While there have been various techniques like question-answer generation as a precursor task to label prediction as in the AVeriTeC shared task, all these shared tasks are centered on predicting a veracity label for a claim. Given the LLM’s tendency to hallucinate generating plausible yet factually inaccurate text, there is a need to organize shared tasks to evaluate the factual accuracy of LLM-generated content.
Open Challenges

### v2

@@ 4.3 Transfer Learning Strategies
Although recent studies show the success of pre-trained LLMs on zero- or few-shot tasks, they often fail at verifying real-world claims given their limited internal knowledge. The success of fine-tuning has motivated recent work on claim verification. Chen et al. Chen et al. (2022b) fine-tuned an LM on an external corpus for retrieving passage titles and evidence sentences using constrained beam search, finding improved performance on the FEVER dataset. Other work demonstrates that using GPT models to generate synthetic training data improves the performance of LLMs on fact checking Tang et al. (2024) and claim matching Choi and Ferrara (2024).
Pan et al. Pan et al. (2021) develop a pipeline for creating a fact verification dataset and fine-tuning a language model by leveraging passages from Wikipedia to generate QA pairs related to claim veracity, showing improved zero-shot performance. Zeng and Zubiaga Zeng and Zubiaga (2024) show that unlabelled pairwise data can increase the alignment between claim-evidence pairs, resulting in significant improvement on few-shot claim verification. Other recent works leverage reinforcement learning to fine-tune models for improving the veracity of claims and supporting evidence Zhang and Gao (2024); Huang et al. (2024). A document-level and question-level retrieval policy is proposed by Zhang and Gao Zhang and Gao (2024), where the top-k and top-1 documents for the document and question-level policy respectively are used as input to a scoring function for label prediction during training. This approach outperforms retrieval and prompting approaches. Chiang et al. Chiang et al. (2024) fine-tune LLMs for multi-stage fact verification. They fine-tune a model to generate answers based on claim-evidence pairs and a set of questions, whereas another model is fine-tuned to verify the claim based on the claim-evidence and question-answer pairs. Zhu et al. Zhu et al. (2023) fine-tune a generation model to generate counterfactuals for out-of-domain classification.
@@
The recent success of pre-trained LLMs in zero- and few-shot settings is largely attributed to its in-context learning (ICL) abilities Kojima et al. (2022); Brown et al. (2020). For claim verification, popular ICL techniques include chain-of-thought (CoT) reasoning Wei et al. (2022). Zhao et al. Zhao et al. (2024) develop a multi-stage verification pipeline based on claim decomposition and self-reflection. An LLM-based verifier module is created using instruction prompting to generate a reasoning analysis among all sub-claims created by the decomposer module. The suggests that zero-shot prompting techniques result in better multi-hop performance on the HOVER and FEVEROUS datasets compared to few-shot prompting and fine-tuning. Kanaani Kanaani (2024) enable LLMs to generate reasons over retrieved evidence in claim verification using few-shot ICL and the STaR CoT technique inspired by Zelikman et al. Zelikman et al. (2022). Similar work has leveraged CoT for effectively verifying complex claims using reasoning steps Yao et al. (2023); Ni et al. (2024). Conversely, HiSS Zhang and Gao (2023) demonstrates that prompting for few-shot learning and claim decomposition can substantially improve the performance of CoT models for complex news claim verification. Li et al. Li et al. (2023b) leverage the ICL capability of LLMs to perform multiple tasks simultaneously. Their approach outperforms or achieves comparable task performance in a zero-shot setting on claim verification datasets.
4.4 LLM Generation Strategies
@@
While the majority of claim verification systems predict veracity based on the concatenation of the input claim and evidence sentences Pradeep et al. (2021), recent work has proposed alternate strategies for determining veracity labels and selecting/generating evidence pieces. Cao et al. Cao et al. (2024) develop SERIf, a claim verification pipeline that features an inference module to predict the veracity label of scientific news articles based on a two-step summarization (i.e., ‘Extractive - Abstractive‘) and evidence retrieval technique. Each summary-evidence pair is fed into the LLM and produces a binary label, indicating whether the news article is reliable (supported) or unreliable (refuted). Wadden et al. Wadden et al. (2022b) leverage the Longformer model Beltagy et al. (2020) that uses a shared encoding over the claim and document abstracts for rationale identification and claim label prediction.
@@
Li et al. Li et al. (2024a) propose to select the minimal evidence group within a set of retrieved candidate documents. This aims to minimize redundancy while selecting the most relevant evidence to prompt the language model. Chen et al. Chen et al. (2022b) use BART to encode all candidate sentences from the most relevant retrieved documents. In this, BART serves as an evidence decoder to predict the $g$-th evidence sentence via generation conditioned on the top $k$ retrieved documents and the input claim. Lee et al. Lee et al. (2022) develop a variant of nucleus sampling called factual-nucleus sampling, in which the top-$p$ sampling pool is selected as a set of sub-words whose cumulative probability exceeds $p$, resulting in improved evidence and label generation without claim verification datasets such as FEVER. Kao and Yen Kao and Yen (2024) propose a multi-stage approach, where the evidence sentences are retrieved from articles related to a claim, and arguments are generated by aggregating and reconstructing the evidence. The arguments are refined and passed to a LLM to generate a verification label. Other approaches leverage LLMs reasoning capabilities to generate veracity labels and factual evidence Cheng et al. (2024); Jafari and Allan (2024); Li et al. (2024b); Fang et al. (2024); Pan et al. (2023a).
@@
Recent studies investigate explainable approaches to improve LLM-based claim verification. Wang and Shu Wang and Shu (2023) present the FOLK framework, that leverages the explanation capabilities of LLMs when verifying claims and justifies the prediction through a summary of its decision process. Dammu et al. Dammu et al. (2024) proposes a knowledge graph (KG)-based approach for text verification and evidence attribution, where the LLM is fine-tuned on evidence attribution based on the input text and retrieved triplets from KG, inducing explanations for claim predictions. While explanation techniques can help humans verify facts, LLMs can produce incorrect explanations due to hallucinations, making them unreliable in certain claim verification scenarios Si et al. (2023). Ma et al. Ma et al. (2024) prompt an LLM in a few-shot setting to generate a concise summary of the evidence documents and input claim, serving as an explanation for the verified claim. Other works leverage reasoning techniques such as chain-of-thought (CoT) to enable the LLM to be interpretable in its decision-making process when verifying claims Yao et al. (2023); Pan et al. (2023a); Zhao et al. (2024); Kanaani (2024); Ni et al. (2024); Quelle and Bovet (2024); Fang et al. (2024). Pan et al. Pan et al. (2023a) and Fang et al. Fang et al. (2024) leverage an LLM’s reasoning ability to generate explanations by using question-guided reasoning and minimizing the inherent model biases.
5 Evaluation and Benchmarking
5.1 Metrics
@@
The F1 score is the most commonly used metric to measure the performance of automatic claim verification systems, and to a lesser degree Precision, Recall, and Accuracy. Katranidis and Barany Katranidis and Barany (2024) used the error rate between the human and automated fact-verification system to measure verification accuracy. However, these metrics consider a single pipeline component to evaluate the system’s overall performance. Hence, Thorne et al. Thorne et al. (2018a) introduce the FEVER score, a metric that incorporates both verification accuracy and evidence retrieval accuracy to compute overall system performance. While these metrics are valuable for assessing the performance of the classification tasks, they are inadequate for evaluating the performance of the non-classification components of the pipeline. Hence, metrics like Recall@k are used to measure the performance of the retrieval task Pan et al. (2023b); Pradeep et al. (2021), while BLEU and METEOR are used to evaluate the quality of explanations or generated questions and answers. Schlichtkrull et al. Schlichtkrull et al. (2024) propose the new evaluation metric AVeriTeC score that uses METEOR and accuracy for question-answer-based veracity prediction systems. Other metrics like Mean Absolute Error (MAE), Expected Calibration Error (ECE), Area Under ROC Curve (AUC-ROC), and Pearson’s Correlation are also used. Most of these metrics are inadequate for evaluating the factual accuracy of LLM-generated text, for which FactScore Min et al. (2023), SAFE Wei et al. (2024), and VERISCORE Song et al. (2024) have been proposed instead. Other metrics and frameworks evaluate factual errors in generated text Lee et al. (2022); Chern et al. (2023) and their alignment Zha et al. (2023) and entailment Lee et al. (2022) considering factuality.
5.2 Datasets
@@
The fundamental resource for training and evaluating claim verification systems is datasets containing annotated texts. As most research in this area deals with English data, we collect information about all publicly available English datasets used in the papers discussed in this survey [^4] https://github.com/LanguageTechnologyLab/Claim-Verification-Papers.git.
@@
General-domain datasets have been created from online data sources and websites including Wikipedia Thorne et al. (2018a); Jiang et al. (2020); Diggelmann et al. (2020); Eisenschlos et al. (2021); Schuster et al. (2021); Kamoi et al. (2023), due to the extensive amount of information available spanning various topics and domains, and online fact-checking websites like PolitiFact Wang (2017); Augenstein et al. (2019); Kao and Yen (2024). Factually incorrect claims are shared through social media channels like X, Reddit, etc. Saakyan et al. Saakyan et al. (2021) introduce the COVID-Fact dataset consisting of claims extracted from Reddit posts. Several datasets for scientific fact verification have also been introduced Wadden et al. (2020, 2022a); Lu et al. (2023). Most fact-verification datasets rely on unstructured textual evidence. Hence, a few datasets with structured evidence sources have been introduced Aly et al. (2021); Lu et al. (2023). While many datasets focus only on veracity labels, some were developed to address explainability too Chen et al. (2022d); Yang et al. (2022); Ma et al. (2024); Rani et al. (2023); Schlichtkrull et al. (2024).
@@
Most claim-verification datasets include claims extracted from available information sources. LLMs are widely used to generate content, even though they can generate factually incorrect information. Efforts to identify such non-factual text have been undertaken. However, claim-verification datasets consisting of human-written claims can be inadequate for this task due to the linguistic discrepancy between human- and LLM-generated text. As such, LLM-generated claim verification datasets have been introduced Li et al. (2023c); Cao et al. (2024). While these are used to evaluate automatic claim verification systems, there is a need to evaluate the factual accuracy of LLM-generated text, leading to the introduction of evaluation datasets Lee et al. (2022); Wang et al. (2024); Malaviya et al. (2024).
5.3 Shared Tasks
@@
Shared tasks are competitions where teams develop systems for tasks using a common benchmark dataset. Multiple shared tasks on claim and fact-checking have been organized including the Fact Extraction and Verification (FEVER) challenge Thorne et al. (2018b), TabFact [^5] https://competitions.codalab.org/competitions/21611, CLEF 2020 CheckThat! Barrón-Cedeno et al. (2020), SCIVER Wadden and Lo (2021), SEM-TAB-FACT Wang et al. (2021), FACTIFY-5WQA [^6] https://defactify.com/factify3.html, and more recently AVeriTeC [^7] https://fever.ai/task.html. While there have been various techniques like question-answer generation as a precursor to label prediction as in AVeriTeC, all these shared tasks are centered on predicting veracity labels. Given an LLM’s tendency to generate plausible yet factually inaccurate text, there is a need to organize shared tasks to evaluate the factual accuracy of LLM-generated content.
6 Open Challenges